# GraphRAG-R1: Graph Retrieval-Augmented Generation with Process-Constrained Reinforcement Learning

### v1

@@ 3.1.1. Rollout-with-Thinking
The model is guided to utilize the external retrieval tool during generation by using two special tokens <begin_of_query>...<end_of_query> to indicate the call to the search tool. When <|end_of_query|> appears, it signifies that the retrieval query has been fully generated. At this point, model generation is paused, the external retrieval tool is called, and the retrieved content is marked with <|begin_of_documents|> ... <|end_of_documents|>. This content then guides the model’s next step of reasoning. This method ensures retrieval is seamlessly integrated into the reasoning process, allowing the model to continue its reasoning based on the retrieved documents without interruption. Detailed prompts are provided in Appendix A.4
@@ 3.1.2. Retrieval-Masked Loss
In the original GRPO, token-level loss is computed throughout the rollout sequence. In the GraphRAG scenario, the rollout sequence consists of tokens generated by the LLM and tokens retrieved from external channels. Applying the same optimization to retrieved tokens and LLM-generated tokens may negatively impact the training process. Therefore, <begin_of_documents> and <end_of_documents> are designated as special tokens and their enclosed content is masked during training, ensuring the policy gradient objective is calculated exclusively on LLM-generated tokens. This approach stabilizes training while preserving the flexibility of retrieval-augmented generation.
@@ 3.2.1. Format reward
This reward function is designed to steer the model toward strictly adhering to the specified output format, ensuring that it learns to correctly invoke external retrieval and generate the final answer. The model should place the final response entirely within the <answer>...</answer> tags and guarantee that the entire output is fully readable and free of uninterpretable artifacts. When invoking retrieval, the model should encapsulate retrieval queries within the <|begin_of_query|>...<|end_of_query|> tags with retrieved information contained within <|begin_of_documents|> ... <|end_of_documents|>. The model should explicitly invoke the retrieval mechanism through this query process rather than directly generating content without retrieval. An illustration of the format is shown in Figure 4 in Appendix A.4.
@@ 3.2.2. Progressive Retrieval Attenuation (PRA)
[…]
@@
Through this reward function, we can effectively restrict the number of model-triggered retrieval operations, effectively preventing shallow retrieval. For different data characteristics, the hyperparameter $k$ can be flexibly adjusted to better steer the model’s learning process.
@@ 3.2.3. Cost-Aware F1 (CAF)
Although the PRA reward can regulate the retrieval frequency of the model within a reasonable range, this mechanism alone still relies on prior knowledge and could not fully activate the reasoning capabilities of LLMs. For example, the LLM may adopt unnecessarily numerous of retrievals, leading to the “over-thinking” problem and greatly waste the computational resources. An ideal GraphRAG model should deliver high-quality responses with precisely needed retrieval operations, which requires three core competencies: accurately assessing the problem difficulty, intelligently decomposing complex queries, and dynamically optimizing retrieval strategies.
@@
where $F1$ is the F1 score of the generated answers, $a$ and $b$ are hyperparameters to control the coefficients in the regulatory function, and $N$ denotes the total number of retrieval operations. The CAF rewards ensure that responses attaining identical F1 scores receive higher rewards when using fewer retrievals, thereby incentivizing the model to proactively recognize problem complexity and refine retrieval decisions. Consequently, CAF reward establishes a balance among reasoning depth, retrieval cost, and generation quality, effectively mitigating suboptimal behaviors of over-thinking.
@@ 3.2.4. Phase-Dependent Training Strategy
In the second stage, we aim to shape the behavior of the model, synergistically combining the format reward and the PRA reward. This design recognizes standardized tool-calling protocols and retrieval behavior as essential pre-requisites for high-quality reasoning. By learning with two objectives, the model can efficiently learn how and when to invoke the retrieval tools with structured output patterns. The decay mechanism of PRA dynamically regulates retrieval behavior through a high initial reward, driving essential information retrieval, while its exponential decay characteristic suppresses redundant operations.
@@
In summary, our phase-dependent training strategy resolves potential gradient conflicts, improving training stability and significantly boosting model performance. The pseudo-code of the training process is shown in Appendix B.1.
@@ 3.3. Hybrid Graph-Textual Retrieval
By employing hybrid retrieval, our method has several advantages. Compared to lengthy text fragments, the compact nature of graph can drastically reduce token consumption, improving processing efficiency. This structured representation offers superior information density, allowing the model to focus directly on essential entity-relationships while inherently filtering out the redundant noise prevalent in raw text snippets. Therefore, the integration of structured knowledge significantly enhances the quality and efficiency of generated answers, as the reasoning operates on a cleaner and more relevant information foundation. We empirically verify the hybrid graph-textual retrieval in Section 4.3.3.
@@
[…]
@@ 4.1.2. Metrics
Following the common literature, we employ F1 Score (Wang et al., 2024), SBERT Similarity (SBERT) (Thakur et al., [n. d.]), and LLM-as-Judge Accuracy ($\mathrm{ACC_{L}}$)  (Song et al., 2025) as evaluation metrics. The F1 Score measures the lexical overlap between generated answers and ground-truth references. SBERT uses the cosine similarity between sentence embeddings generated by the SBERT model to measure semantic similarity. $\mathrm{ACC_{L}}$ leverages an LLM (specifically, Qwen-3-Turbo) as an evaluator to assess the correctness of predicted answers based on contextual understanding.
@@ 4.1.3. Baselines
We compare 7 representative GraphRAG methods:
  • KGP (Wang et al., 2024): It utilizes fine-tuned LLMs to explore neighboring nodes and aggregate information from a KG to produce answer.
@@ 4.1.4. Implementation Details
We utilize Qwen-2.5-7B as the default backbone LLM model. For baselines that require external closed-source LLMs, we utilize Qwen-3-Turbo. We use HippoRAG2 as the external retriever for LLM’s rollout retrieval.
[…]
@@ 4.2. Main Results
GraphRAG-R1  achieves consistent improvements across all datasets. Specifically, we achieve the best results for evaluation metrics F1 and SBERT on all datasets, and achieve three best results and one second-best result for $\mathrm{ACC_{L}}$. This fully validates the effectiveness of the GraphRAG-R1  framework in enhancing LLMs’ capability to handle multi-hop reasoning problems. Particularly, it outperforms the strongest baseline in F1 score by up to 38.03% on HotpotQA, 62.43% on MuSiQue, and 83.81% on 2Wiki, respectively. GraphRAG-R1  get more improvement on a more difficult multi-hop dataset. This demonstrates the framework’s efficacy in enhancing large language models’ multi-hop reasoning capabilities.
GraphRAG-R1 achieves leading performance on both in-domain datasets (HotpotQA, MuSiQue, and 2Wiki) and the out-of-domain dataset PopQA. Concretely, it outperforms the best baseline by up to 19.96% in terms of F1 on PopQA, indicating that our proposed method possesses strong generalization ability and can effectively handle unseen complex questions. In comparison, PropRAG, which show the most competitive performance on HotpotQA and MuSiQue, do not generalize to the out-of-domain dataset, indicating generalization is challenging for existing GraphRAG methods.
From the results of self-constructed baselines, existing techniques such as naive RAG, prompt engineering, and SFT can indeed improve the ability of vanilla LLM on complex question answering tasks. However, our method, which adopts RL-based training, is clearly more effective, outperforming all these non-RL approaches usually by a large margin. The consistent performance improvement validates that RL-based methods hold more potential in enhancing GraphRAG and the reasoning ability of LLMs in general.
@@
[…]
@@ 4.3.1. PRA & CAF Rewards
[…]
@@
Besides the quantitative F1 score, we also delve deeper into the results by reporting the number of retrieval calls. We make the following observations. Without using PRA, our method has a reduced number of retrieval calls, indicating that it suffers from the “shallow retrieve” problem and leads to unsatisfactory performance. PRA dynamically regulates and encourages retrieval calls through its decay factor, making the retrieval behavior more controllable.
@@
[…]
@@
On the other hand, without using CAF, the number of retrieval calls consistently increases across all datasets, indicating the model is more inclined to invoke external functions. However, this not yield to corresponding F1 improvements in answer quality. In fact, the F1 score is even lower than methods invoking few retrievals, indicating that the model tends towards “over-thinking” and performing unnecessary retrievals, potentially due to reward-hacking of solely using the PRA reward.
When both rewards are adopted, our method achieves the best performance with a suitable number of retrieval calls, indicating our method can calibrate its retrieval strategy to the inherent difficulty of the problem. For HotpotQA and PopQA, which contain relatively simpler problems, the number of retrieval calls is automatically decreased. Conversely, for inherently more complex problems demanding deeper information gathering, such as MuSiQue and 2Wiki, our model learns to invoke retrievals more frequently. In sum, our method can dynamically balance efficiency and effectiveness through reward designs.
@@ 4.3.2. Phase-Dependent Training Strategy
To validate the effectiveness of our training strategy, we compare with two variants: one without using the first cold-start stage (denoted as “-cold start”), and the other one without using stages, i.e., directly mixing all reward functions for training (denoted as “-all stages”). The results are shown in Table 3. We have the following observations.
@@
Compared to training, which used format, PRA, and CAF rewards simultaneously, we found that the three-stage training strategy delivers superior performance. The first stage, which focused on format and tool invocation learning, establishes a more solid foundation for optimizing answer quality in the second stage. This approach avoids conflicts or convergence difficulties caused by overly complex reward signals.
@@ 4.3.3. Hybrid Graph-Textual Retrieval
To demonstrate the effectiveness of our hybrid graph-textual retrieval design, we compare different retrieval strategies. During training, we adopt three settings: using mixing text and graph (denoted as “Text+Graph”), text retrieval only (denoted as “Text”), and using graph only (denoted as ”Graph”). During inference, we utilize 4 different retrieval methods: using 5 text snippets (denoted as “Text5”), using 3 text snippets and 10 graph triplets (denoted as “Text3+Graph10”), using 1 text snippet and 10 graph triplets (denoted as “Text1+Graph10”), and using 20 graph triplets(denoted as “Graph20”). The results are shown in Figure 3 and we make the following two observations.
@@
For different inference strategies, as the proportion of graph data increases, the number of tokens consumed significantly reduces, while the F1 score of the model only slightly decreases. This demonstrates that graph information generally possess high information density, enabling the capture of valuable information needed for question answering. Our hybrid retrieval approach can achieve a suitable balance between model effectiveness and computational resource consumption.
@@ 4.3.4. Integration with Other Retrievers
To answer Q3, we conduct experiments to integrate our proposed method with other existing retrieval methods. Specifically, we choose all RAG methods used in our baselines, including Naive RAG, ToG, KGP, LightRAG, PropRAG, G-retriever, and HippoRAG2. As shown in Table 4, GraphRAG-R1 generally improves the performance of different RAG methods across datasets. Specifically, GraphRAG-R1 achieves notable average improvements over the original methods, with the highest enhancement of 38.37% on LightRAG and consistent gains ranging from 3.52% to 36.25% in other baselines. Note that our method is only trained once and only has access to HippoRAG2 during training, indicating that the reasoning and retrieval ability learned by GraphRAG-R1 can be easily transferred and generalized to other GraphRAG methods.
More experiments including adopting LLaMA-3 as the backbone LLM, a comparison of the number of tokens, and detailed prompts used are provided in the Appendix.
@@ Appendix A More Experiments
A.1. More Details about Datasets

### v2

@@ 3.1.1. Rollout-with-Thinking
The model is guided to utilize the external retrieval tool during generation by using two special tokens <|begin_of_query|>...<|end_of_query|> to indicate the call to the search tool. When <|end_of_query|> appears, it signifies that the retrieval query has been fully generated. At this point, model generation is paused, the external retrieval tool is called, and the retrieved content is marked with <|begin_of_documents|> ... <|end_of_documents|>. This content then guides the model’s next step of reasoning. This method ensures retrieval is seamlessly integrated into the reasoning process.
@@ 3.1.2. Retrieval-Masked Loss
In the original GRPO, token-level loss is computed throughout the rollout sequence. In the GraphRAG scenario, the rollout sequence consists of tokens generated by the LLM and tokens retrieved from external channels. Applying the same optimization to retrieved tokens and LLM-generated tokens may negatively impact the training process. Therefore, we utilize special tokens <|begin_of_documents|> and <|end_of_documents|> and their enclosed content is masked during training, ensuring the policy gradient objective is calculated only on LLM-generated tokens.
@@ 3.2.1. Format reward
This reward function is designed to steer the model toward strictly adhering to the specified output format, ensuring that it learns to correctly invoke external retrieval and generate the final answer. The model should place the final response entirely within the <answer>...</answer> tags and ensure that the entire output is fully readable and free of uninterpretable artifacts. When invoking retrieval, the model should encapsulate retrieval queries within the <|begin_of_query|>...<|end_of_query|> tags with retrieved information contained within <|begin_of_documents|> ... <|end_of_documents|>. The model should explicitly invoke the retrieval mechanism through this query process rather than directly generating content without retrieval.
@@ 3.2.2. Progressive Retrieval Attenuation (PRA)
where $R_{0}$ is the base reward, $n$ is the number of retrieval calls, and $0<k<1$ is the decay factor. Through this dynamically adjusted reward weighting, the retrieval frequency of the model can be precisely controlled within the desired threshold ranges. Using the properties of geometric series, it is easy to see that:
@@
Through this reward function, we can effectively restrict the number of retrieval calls, effectively preventing shallow retrieval. For different data characteristics, the hyperparameter $k$ can be flexibly adjusted to better steer the model’s learning process.
@@ 3.2.3. Cost-Aware F1 (CAF)
Although the PRA reward can regulate the retrieval frequency of the model within a reasonable range, this mechanism alone still relies on prior knowledge and could not fully activate the reasoning capabilities of LLMs. For example, the LLM may adopt unnecessarily numerous retrievals, leading to the “over-thinking” problem and greatly waste the computational resources. An ideal GraphRAG model should deliver high-quality responses with precisely needed retrieval operations, which requires three core competencies: accurately assessing the problem difficulty, intelligently decomposing complex queries, and dynamically optimizing retrieval strategies.
@@
where $F1$ is the F1 score of the generated answers, $a$ and $b$ are hyperparameters to control the coefficients in the regulatory function, and $N$ denotes the total number of retrieval operations. The CAF reward ensures that responses achieving identical F1 score receive higher rewards when using fewer retrievals, thereby incentivizing the model to proactively recognize problem complexity and refine retrieval decisions. Consequently, CAF reward establishes a balance among reasoning depth, retrieval cost, and generation quality, effectively mitigating suboptimal behaviors of over-thinking.
@@ 3.2.4. Phase-Dependent Training Strategy
In the second stage, we aim to shape the behavior of the model, synergistically combining the format reward and the PRA reward. This design recognizes standardized tool-calling protocols and retrieval behavior as essential prerequisites for high-quality reasoning. By learning with two objectives, the model can efficiently learn how and when to invoke the retrieval tools with structured output patterns. The decay mechanism of PRA dynamically regulates retrieval behavior through a high initial reward, driving essential information retrieval, while its exponential decay characteristic suppresses redundant operations.
@@
In summary, our phase-dependent training strategy resolves potential gradient conflicts, improving training stability and significantly boosting model performance.
@@ 3.3. Hybrid Graph-Textual Retrieval
By employing hybrid retrieval, our method has several advantages. Compared to lengthy text fragments, the compact nature of graph can drastically reduce token consumption, improving processing efficiency. This structured representation allows the model to focus directly on essential entity-relationships while inherently filtering out the redundant noise. Therefore, the integration of structured knowledge significantly enhances the quality and efficiency of generated answers, as reasoning operates on a cleaner and more relevant information foundation. We empirically verify the hybrid graph-textual retrieval in Section 4.3.3.
@@
$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
@@ 4.1.2. Metrics
Following the common literature, we employ F1 Score (Wang et al., 2024), SBERT Similarity (SBERT) (Thakur et al., [n. d.]), and LLM-as-Judge Accuracy ($\mathrm{ACC_{L}}$)  (Song et al., 2025) as evaluation metrics. The F1 Score measures the lexical overlap between generated answers and ground-truth references. SBERT uses the cosine similarity between sentence embeddings generated by the SBERT model to measure semantic similarity. $\mathrm{ACC_{L}}$ leverages an LLM (specifically, Qwen3-Turbo) as an evaluator.
@@ 4.1.3. Baselines
We compare 7 representative methods:
  • KGP (Wang et al., 2024): It utilizes fine-tuned LLMs to explore neighboring nodes and aggregate information from a KG to produce answers.
@@ 4.1.4. Implementation Details
We utilize Qwen2.5-7B as the default backbone LLM model. For baselines that require external closed-source LLMs, we use Qwen3-Turbo. We use HippoRAG2 as the external retriever for LLM’s rollout retrieval.
For training, in the first stage, we generate training data using ERNIE-3.5-128K (Wang et al., 2021). In the second stage, we set the base reward for PRA as $R_{0}=0.5$ and its decay factor as $k=0.5$. In the third stage, we set the hyper-parameter for CAF as $a=2$ and $b=0.1$. Throughout three stages, LoRA fine-tuning (Hu et al., 2021) is employed with $r=16$ and $\text{lora}\_{\text{alpha}}=32$. The sampling temperature is set to $1.0$ with a maximum retrieval count of $8$, the learning rate is 5e65e-6, and the number of epochs for SFT is $1$.
@@ 4.2. Main Results
GraphRAG-R1  achieves consistent improvements across all datasets. Specifically, we achieve the best results for evaluation metrics F1 and SBERT in all datasets, and achieve three best results and one third-best result for $\mathrm{ACC_{L}}$. This fully validates the effectiveness of the GraphRAG-R1  framework in enhancing LLMs’ capability to handle multi-hop reasoning problems. Particularly, it outperforms the strongest baseline in F1 score by up to 38.08% on HotpotQA, 62.43% on MuSiQue, and 83.81% on 2Wiki, respectively. GraphRAG-R1  gets more improvement on a more difficult multi-hop dataset.
GraphRAG-R1 achieves leading performance on both in-domain datasets and the out-of-domain dataset PopQA. Concretely, it outperforms the best baseline by up to 19.96% in terms of F1 on PopQA, indicating that our proposed method possesses strong generalization ability and can effectively handle unseen complex questions.
From the results of self-constructed baselines, existing techniques such as naive RAG, prompt engineering, and SFT can indeed improve the ability of vanilla LLM on complex question answering tasks. However, our method, which adopts RL-based training, is clearly more effective, usually outperforming all these non-RL approaches by a large margin. The consistent performance improvement validates that RL-based methods hold more potential in enhancing GraphRAG and the reasoning ability of LLMs in general.
@@
$\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$
@@ 4.3.1. PRA & CAF Rewards
$\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$
@@
Besides the quantitative F1 score, we also delve deeper into the results by reporting the number of retrieval calls. We make the following observations. Without using PRA, our method has a reduced number of retrieval calls, indicating that it suffers from the “shallow retrieval” problem and leads to unsatisfactory performance. PRA makes the retrieval behavior more controllable.
@@
$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
@@
On the other hand, without using CAF, the number of retrieval calls consistently increases across all datasets, indicating the model is more inclined to invoke external functions. However, this not yield to corresponding F1 improvements in answer quality. In fact, the F1 score is even lower than methods invoking few retrievals, indicating that the model tends towards “over-thinking” and performing unnecessary retrievals.
When both rewards are adopted, our method achieves the best performance with a suitable number of retrieval calls, indicating our method can calibrate its retrieval strategy to the inherent difficulty of the problem. For HotpotQA and PopQA, which contain relatively simpler problems, the number of retrieval calls is automatically decreased. Conversely, for inherently more complex problems demanding deeper information gathering, such as MuSiQue and 2Wiki, our model learns to invoke retrievals more frequently.
@@ 4.3.2. Phase-Dependent Training Strategy
To validate the effectiveness of our training strategy, we compare our method with two variants: one without the first cold-start stage (denoted as “-cold start”), and the other without distinct stages, i.e., directly mixing all reward functions for training (denoted as “-all stages”). The results are shown in Table  3. Here, $\mathrm{ACC_{F}}$ represents the percentage of model outputs that are entirely accurate in format.
@@
Compared to training, which used format, PRA, and CAF rewards simultaneously, we found that the three-stage training strategy delivers superior performance. This approach avoids conflicts or convergence difficulties caused by overly complex reward signals.
@@ 4.3.3. Hybrid Graph-Textual Retrieval
To demonstrate the effectiveness of our hybrid graph-textual retrieval design, we compare different retrieval strategies. During training, we adopt three settings: using mixing text and graph (“Text+Graph”), text retrieval only (“Text”), and using graph only (”Graph”). During inference, we utilize 4 different retrieval methods: using 5 text snippets (“Text5”), using 3 text snippets and 10 graph triplets (“Text3+Graph10”), using 1 text snippet and 10 graph triplets (“Text1+Graph10”), and using 20 graph triplets ( “Graph20”). The results are shown in Figure  3.
@@
For different inference strategies, as the proportion of graph data increases, the number of tokens consumed significantly reduces, while the F1 score of the model only slightly decreases. This demonstrates that our hybrid retrieval approach can achieve a suitable balance between model effectiveness and computational resource consumption.
@@ 4.3.4. Integration with Other Retrievers
To answer Q3, we conduct experiments to integrate our proposed method with other existing retrieval methods. As shown in Table 4, GraphRAG-R1 generally improves the performance of different RAG methods across datasets. The last column ”Imp” is defined as the average percentage increase in all metrics after integrating GraphRAG-R1  relative to original method. Specifically, GraphRAG-R1 achieves notable average improvements over the original methods, with the highest improvement of 38.37% on LightRAG and consistent gains ranging from 5.26% to 36.25% in other baselines. Note that our method is only trained once and only has access to HippoRAG2 during training, indicating that the reasoning and retrieval ability learned by GraphRAG-R1 can be easily generalized to other GraphRAG methods. More experiments are provided in the Appendix.
@@ Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (No. 92570113, 62472018), the Natural Science Foundation of Tianjin (No. 24JCQNJC02170), and the Tianjin Science and Technology Plan Project (No. 24HHXCSS00001).
@@ Appendix A More Experiments
A.1. Experiments with Qwen2.5-7B-Instruct
In this study, we further explore the performance of GraphRAG-R1 on instruct-tuned models. We observed that results based on the original Qwen2.5-7B-base still have room for improvement on certain tasks, particularly in terms of instruction compliance and contextual understanding. Therefore, we selected the Qwen2.5-7B-Instruct model as the backbone. Compared to the base model, the instruct-tuned model demonstrates stronger capabilities in understanding user intent, following output formats, and handling multi-turn dialogues. Based on this model, we retained the original GraphRAG-R1 framework and only performed adaptive tuning on the training strategy, prompt templates, and inference hyperparameters, aiming to verify the stability and scalability of our method across different types of backbone models.
The experimental results demonstrate that the tuned GraphRAG-R1 with Qwen2.5-7B-Instruct not only significantly outperforms the base version but also achieves the best performance on most evaluation metrics across datasets such as HotpotQA, MuSiQue, 2Wiki, and PopQA. Detailed performance comparisons are provided in Table 5.
Furthermore, we tested the integration of our approach with other retrievers, as shown in Table 6. In all tested configurations, GraphRAG-R1 consistently improved performance over the original retrieval methods, demonstrating strong transferability and generalization. This indicates that our framework can be flexibly adapted to different retrieval mechanisms.
A.2. Experiments with LLaMA-3-8B
Table 7. The results of different methods for multi-hop question answering using Llama-3-8B as the backbone LLM. The best results are in bold and the second-best results are underlined.
Method HotpotQA MuSiQue 2Wiki PopQA
$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
Naive RAG 9.57 27.00 48.25 1.17 4.00 40.11 4.14 10.00 43.58 8.14 24.70 46.59
KGP 3.61 17.00 46.41 1.38 3.50 44.49 8.23 26.50 48.48 10.75 59.00 48.13
ToG 8.97 16.50 49.83 3.28 10.00 44.74 7.91 23.00 48.36 10.96 37.50 49.57
LightRAG 5.33 12.50 43.95 1.83 3.50 41.51 4.68 12.00 43.95 6.71 27.50 45.83
PropRAG 12.47 23.00 49.03 3.83 6.50 42.17 5.06 11.50 43.06 9.76 25.70 48.10
G-retriever 1.91 23.00 43.09 1.06 12.50 42.77 2.43 18.00 42.26 0.25 38.10 39.32
HippoRAG2 6.89 22.00 46.06 1.85 4.00 42.03 5.61 12.00 43.51 7.85 26.10 46.57
GraphRAG-R1 13.16 26.50 53.06 6.32 16.00 48.96 13.36 25.00 51.46 11.64 41.50 50.42
In this subsection, we investigate the effectiveness of GraphRAG-R1  across different base models. We adopt another representative open-sourced LLM, Llama-3-8B, as the backbone LLM. The experimental setting is similar to using Qwen2.5-7B, except we add PRA reward in the third training stage to encourage the model to invoke retrieval, which enhances performance. The overall results are presented in Table 7. Collectively, these experiments demonstrate that our method is effective and outperforms baselines when using LLaMA-3-8B as the backbone LLM, showing that our method is generalizable across different LLM models.
A.3. Hyperparameter Sensitivity Analysis
To evaluate the hyperparameter robustness of the GraphRAG-R1 framework, we conducted sensitivity analysis on four key parameters. The results are shown in Figure 4 and demonstrate that the performance of GraphRAG-R1 has a wide robustness interval around the parameters selected in the main experiment, which enhances the reproducibility and practical reliability of the method.
Figure 4. Sensitivity analysis of key hyperparameters in GraphRAG-R1. Subfigures (a)-(d) illustrate the variation in F1 score on the four validation sets as the key hyperparameters are varied within specified ranges. Other hyperparameters were fixed at their main experiment values. The gray dashed lines indicate the parameter values selected for the main experiments.
Sensitivity analysis of key hyperparameters in GraphRAG-R1. Subfigures (a)-(d) illustrate the variation in F1 score on the four validation sets as the key hyperparameters are varied within specified ranges. Other hyperparameters were fixed at their main experiment values. The gray dashed lines indicate the parameter values selected for the main experiments.
A.4. Cost and Efficiency Analysis
Table 8. Performance and cost comparison with KGP and ToG, two reasoning-enhanced GraphRAG methods
Method HotpotQA MuSiQue 2Wiki PopQA
$\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$
KGP 10.73 2687.44 4.61 2858.23 10.16 2173.02 21.01 3182.02
ToG 11.44 3136.61 5.02 2873.56 14.45 2065.94 29.21 2838.49
GraphRAG-R1 38.00 1674.47 20.06 2086.10 32.24 1883.21 35.04 1449.80
Method Model HotpotQA MuSiQue 2Wiki PopQA Imp
$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
Dalk Original 0.62 7.50 41.96 0.40 3.00 40.64 0.75 4.00 41.66 0.14 8.80 39.37
+GraphRAG-R1 15.95 20.50 52.38 8.15 8.50 48.33 10.45 14.40 49.54 13.38 26.60 51.90 + 69.50%
In this subsection, we evaluate the performance and cost of different reasoning-enhanced retrieval methods. We compare our approach with KGP and ToG, two reasoning-enhanced retrieval methods that also support multiple retrievals. In addition to comparing the performance measured by F1 scores, we also calculate the token consumption throughout the retrieval and reasoning processes. Specifically, for ToG and KGP, we count the number of accesses to the knowledge graph, for our method, we tally the frequency of calls to external retrieval tools. Token overhead accounts for consumption throughout the retrieval and reasoning processes. As shown in Table  8, GraphRAG-R1  achieves significant improvements in both performance and computational cost by deeply integrating the intrinsic reasoning capabilities of LLMs. Compared to baseline methods that also leverage LLM reasoning, our method demonstrates substantially improved answer quality while requiring lower token overhead than the other two methods. This validates the effectiveness and cost efficiency of our approach.
A.5. Cross-Domain Adaptation Verification
GraphRAG-R1 is designed as a general framework agnostic to specific graph structures, tasks, and domains. To verify its domain generalization and compatibility, we adapted an advanced retrieval method DALK (Li et al., 2024b) originally designed for the medical domain, integrated it into our framework, and evaluated it on general-domain datasets. We only replacing GraphRAG-R1’s underlying retrieval tool with the adapted DALK retriever. As shown in Table 9, the integrated framework achieved a significant improvement compared to the original DALK method, with a relative F1 score increase of approximately 69.50% on Qwen2.5-7B. This strongly proves that our framework can flexibly compatibly integrate and enhance specialized retrievers from different domains.
A.6. More Details about Datasets
Table 10. Dataset Statistics
Metric MuSiQue 2Wiki HotpotQA PopQA
Graph # Passages 11,656 6,119 9,221 8,676
# Nodes 91,729 42,694 82,157 74,472
# Triplets 107,448 50,671 98,709 139,273
Questions Total 1,000 1,000 1,000 1,000
Training Set 800 800 800 0
Testing Set 200 200 200 1,000